# A Hierarchy of Spectral Gap Certificates for Frustration-Free Spin Systems

### 06 Nov 2024 v1

@@ III.2 Proof of inclusion
meaning that the Knabe, Gosset–Mozgunov, and ”martingale” bounds (eqs. 21 and 31) calculated for a particular size $n$ of the truncation, yield weaker estimates, than the LTI SDP of the same size. Concretely, in SDP terms, we show that the finite size criteria bounds lie in the feasible set of the LTI SDP. The above relation relies on finding a positive semidefinite generating term $q_{n}(\delta)$ for the truncated operators $Q_{n}(\delta)$ corresponding to each of the bounds. Here we show the expressions for martingale bound,
$q_{n}(1-2\eta_{l}):=\eta_{l}(h_{i}+h_{i+1})+\sum_{x=1}^{n-2}\{h_{i},h_{i+x}\},$ (33)
which corresponds to the lower bound on the gap of the blocked Hamiltonian $K$ (as defined in Method 2 (section III.1)) and the expression for the Knabe bound is,
@@
The detailed proofs and the expression for the Gosset–Mozgunov bound are given in Appendix C.
An important implication of eq. 32 is that the LTI-SDP does not necessarily lower bound the gap of the open boundary system, i.e. $\epsilon_{n}$, in the thermodynamic limit. This is because, the formulation of $\Delta_{\mathrm{LTI}}(n)$ only ensures that it lower bounds the gap of a periodic ring of size $m>2n$. This becomes evident by taking the $n\to\infty$ limit on both sides of eq. 32,
@@
IV Numerical benchmarking
In this section we present the results of the numerical application of the method presented in Section II to 1-dimensional spin-chain models. We compared our method with existing finite-size criteria, namely, the Knabe bound (eq. 22) and the bound due to Gosset and Mozgunov (eq. 23).
Because of the inclusions shown in Sec. III, we know that for the same $n$ our method always gives an equal or tighter lower bound on the gap than either of those existing bounds. However, when comparing the methods we should keep in mind that the computational tasks involved are different. While the finite-size methods require the computation of the gap of the $n$-body OBC Hamiltonian—a task with a memory scaling of $\mathcal{O}(d^{n})$—our method optimizes over a positive semidefinite matrix of size $d^{n}\times d^{n}$. The results in this section show that our method outperforms the existing ones even when compared on equal footing (in terms of size of the largest variable in the problem): For the models on which we benchmarked, we observed that $\Delta_{\text{LTI}}(n)\geq\Delta_{\mathrm{Knabe}}(2n)$, and similarly for $\Delta_{\mathrm{GM}}(2n)$.
@@ IV Numerical benchmarking
Before presenting the results we address a technical but important issue: When performing the optimization on a computer the SDP solver will produce a solution which is feasible up to some finite precision. If we wish to regard such a solution as providing a mathematical proof of the existence of a gap, we have to make sure that we have a way to obtain a strictly feasible solution from the one produced by the solver. In our case we obtain a numerical solution for the variable $q_{n}$ which should serve as our certificate of a gap (via equation eq. 7). In practice we obtain $q_{n}\succeq-\epsilon$ where $\epsilon$ is the solver precision. For the LTI SDP one cannot simply add $\epsilon\mathbb{I}$ to $q_{n}$ as this would shift the global energy by an extensive amount and a more refined approach has to be taken. In Appendix E, we explain how we modified the LTI SDP to obtain certified solutions.
IV.1 The AKLT model
@@
The AKLT model was the first example of a gapped isotropic spin chain with integral spin [43]. When the model was initially introduced it was proven to be gapped but an explicit lower bound was not provided [44]. In the same year, Knabe introduced his method and provided a lower bound of $0.24806$ for the gap of the AKLT model. Garcia-Saez et al. [45], using a method based on the rotation symmetry of the Hamiltonian, estimated the AKLT gap to be $0.350$, which is compatible with both exact diagonalization results for finite periodic chains (0.35012(0.35012( 0.35012 for $n=16$ [45]), as well as DMRG calculations for finite open chains ($0.35037$ for $n=100$  [^1] Mingru Yang, private communication). Our method proves a lower bound of $\Delta_{\text{AKLT}}\geq 0.34976$ which was obtained by solving the LTI SDP eq. 11) for $n=6$.
Figure 2 shows the Knabe (eq. 22) and Gosset–Mozgunov (eq. 23) lower bounds, and the lower bounds computed with our method (the solution of eq. 11), all for different values of $n$. In addition the finite systems gaps obtained using exact diagonalization for periodic boundary conditions (PBC) are plotted for $n=8,12$ and $16$ as horizontal lines.
We observe that while all three methods (ours, Knabe, and Gosset–Mozgunov) detect the gap starting from system size $n=4$, the LTI SDP performs significantly better than any of the other methods. This is true even if we compare $\Delta_{\text{LTI}}(n)$ with $\Delta_{\mathrm{Knabe}}(2n)$ and $\Delta_{\mathrm{GM}}(2n)$. Recall that the LTI SDP eq. 11 provides a lower bound on the gap of any periodic system of size $\geq 2n$. Our results for the AKLT model show that the LTI SDP gets very close to this optimal value: for $n=6$ we have $\Delta_{\text{PBC}}(12)-\Delta_{\text{LTI}}(6)=0.00036$.
@@
Figure 2: The gap of the $S=1$ AKLT chain. Lower bounds on the energy gap above the ground state of the infinite AKLT chain obtained with different methods are plotted as a function of the size $n$ of the finite system: The Knabe bound eq. 22, the Gosset–Mozgunov bound eq. 23, and our LTI SDP method eq. 11. In addition, gaps of finite systems with periodic boundary conditions (PBC) obtained with exact diagonalization (ED) are plotted for rings of sizes $n=8,12$, and $16$ by dotted, dashed, and solid horizontal black lines, which are almost indistinguishable (the gaps of the $n=12$ and $n=16$ rings differ in the 6th digit). The LTI SDP significantly outperforms the other two methods and gives a lower bound of $0.34976$ for the AKLT gap.
Figure 3: Gaps of the $\mathbb{Z}_{3}$ family of deformed Potts clock models. For a family of spin chain models parameterized by two parameters $r,s$ (eq. 37) lower bounds on the gap were computed at different values of the parameters using different methods: the Knabe and Gosset–Mozgunov finite size criteria (eq. 22 and eq. 23 respectively), and our LTI SDP method eq. 11. The shaded regions indicate the parameter values for which the finite-size methods detected a gap using exact diagonalization results for systems up to size $n=10$ (green region with dashed border for Knabe and blue region with dash-dotted border for Gosset-Mozgunov). The different markers indicate parameter values for which the LTI SDP detected a gap, with different markers corresponding to different values of $n$ (see legend). At each point $r,s$ the marker corresponds to the lowest value of $n$ for which the LTI SDP detected a gap. Darker marker colors indicate larger gaps (see color bar). Each marker corresponds to the larger of the two gaps resulting from using either the original Hamiltonian or the modified Hamiltonian—in which the interaction term is a projection—as the input for the LTI SDP.
IV.2 Deformed clock models
@@
Due to the fact that the deformation preserves the product-state structure of the ground states of eq. 36, the martingale method can be used to prove that the model is gapped in the whole parameter range [14]: the overlaps between the the ground-state projectors are easy to compute analytically and $\eta$ in eq. 30 is thus shown to decay with the size of the support of the state ($r$ in eq. 30) (see also [47] for a specific treatment of this family of models). This strategy for proving the existence of a gap, however, has the disadvantage that it does not provide us with a concrete lower bound. The exact lower bound proven in this way has a prefactor: the gap of the blocked interaction term $\tilde{h}$ which acts on $2r\gg 1$ spins.
@@
We computed lower bounds on the gap for values of $(r,s)$ in the region $(0,1]\times(0,1]$ using the Knabe method (eq. 22), Gosset–Mozgunov (eq. 23) and our method (eq. 11). Figure 3 shows the regions in which each of the methods detected a gap. In the figure, the green-shaded region with dashed boundary indicates where Knabe’s criterion could detect a gap for system sizes $n=3,\ldots,10$, i.e., where $\max_{n\leq 10}\Delta_{\text{Knabe}}(n)>0$; the blue-shaded region with dash-dotted boundary marks the same for the Gosset–Mozgunov bound. The results obtained with our LTI method (eq. 11) are plotted in Figure 3 as individual data points. For each point on a grid in the $(r,s)$ plane we plotted a marker indicating the lowest $n$ for which a positive gap was detected by the LTI SDP (triangle for $n=3$, square for 4, and pentagram for 5).
@@
IV.3 1D Glauber model
Figure 4: Gaps of the 1D Glauber Hamiltonian family. For a family of spin chain models parameterized by two parameters $\gamma,\delta$ (eq. 42) lower bounds on the gap were computed at different values of the parameters using different methods: the Knabe and Gosset–Mozgunov finite size criteria (eq. 22 and eq. 23 respectively) computed for the blocked Hamiltonian $\tilde{h}$ (eq. 46), and our LTI SDP method eq. 11 computed for the original Hamiltonian. The green shaded region (dashed border) marks the region in which the Knabe method detected a gap ($\gamma\lesssim 0.92$), the blue shaded region (dash-dotted border) marks the region where the Gosset–Mozgunov bound did ($\gamma\lesssim 0.96$). The gaps detected by our LTI method are plotted as individual data points with marker colors corresponding to the size of the gap (see color bar).
@@
|$\ket{\psi}=\frac{1}{\sqrt{Z_{N}}}\sum_{\sigma}e^{-\beta H(\sigma)/2}\ket{\sigma}$ (41)
where $H(\sigma)$ represents a class of classical kinetic Ising models. For this class of Hamiltonians, the state above is found to be representable as matrix product states, thus allowing the construction of a parent Hamiltonian, i.e. a frustration-free Hamiltonian which has the state as its ground state [48]. Here we consider the family of parent Hamiltonians known as the 1D Glauber model [49], which has the local term
@@
i.e., independent of $\gamma$. Note that for calculating finite size criteria bounds, the local gap threshold $\epsilon_{n}$ (Sec. III) corresponds to the gap of the projector Hamiltonian, which is independent of $\delta$.
Finally, to compute the finite size criteria bounds for a 3-local Hamiltonian, we have to adapt the calculation of $\Delta_{\mathrm{Knabe}}$ and $\Delta_{\mathrm{GM}}$ using a two-step blocking procedure. We first block two $\mathrm{spin}-\frac{1}{2}$ particles of the original chain into one $\mathrm{spin}-\frac{3}{2}$ $(d=4)$ particle,
$\{i,i+1\}$$\rightarrow\mathcal{I}.$ (45)
@@
$\tilde{h}_{\mathcal{I},\mathcal{I}+1}$ $=h_{i,i+1}+h_{i+1,i+2},$ (46)
@@ V Conclusion and Outlook
An interesting question we did not address in this paper is the completeness of the LTI hierarchy. Proving that the lower bound produced by our method converges to the infinite-system gap as $n$ is increased is left for future work. It would also be interesting to obtain a bound on the convergence rate to the exact spectral gap in the thermodynamic limit.
@@
Finally, from its formulation, we believe that our method can allow the certification of gaps for systems in higher spatial dimensions. The generalization of our method to such systems is straightforward. The challenge in applying our method in this setting lies solely in the ability to solve large scale SDPs. Yet, from the results we presented for one-dimensional systems it seems reasonable to expect that the system size for which one would need to solve the problem should be significantly smaller than the system sizes involved in proving gaps using finite size criteria in 2D (e.g. [34]).
Acknowledgements.
We thank Milán Rozmán for guiding us through the martingale literature. We thank Mingru Yang for enlightening discussions as well as for providing us with the DMRG result for the AKLT model. We thank Ignacio Cirac for stimulating discussions at the beginning of the project. We thank Jin-Fu Chen for useful discussions about the Glauber model. I.K. and N.S. acknowledge support from the European Union’s Horizon 2020 research and innovation programme through Grant No. 863476 (ERC-CoG SEQUAM), and the FWF Quantum Austria Funding Initiative project “Entanglement Order Parameters” (Grant DOI 10.55776/P36305), funded through the European Union – NextGenerationEU. N.S. also acknowledges support by the the Austrian Science Fund FWF through Grant DOIs 10.55776/COE1 and 10.55776/F71. K.S.R. and J.T. acknowledge the support received from the European Union’s Horizon Europe research and innovation programme through the ERC StG FINE-TEA-SQUAD (Grant No. 101040729). P.E. and J.T. acknowledge the support received by the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme. P.E. acknowledges the support received through the NWO-Quantum Technology programme (Grant No. NGF.1623.23.006). This publication is part of the ‘Quantum Inspire - the Dutch Quantum Computer in the Cloud’ project (with project number [NWA.1292.19.194]) of the NWA research program ‘Research on Routes by Consortia (ORC)’, which is funded by the Netherlands Organization for Scientific Research (NWO). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the funding institutions. Neither of the funding institutions can be held responsible for them. Parts of this work were performed by using the compute resources from the Academic Leiden Interdisciplinary Cluster Environment (ALICE) provided by Leiden University. F.B. acknowledges financial support from the Alexander von Humboldt Foundation.

### 31 Mar 2026 v2

@@ 3.2 Proof of inclusion
meaning that the Knabe, Gosset–Mozgunov, and ”martingale” bounds (eqs. 22 and 33) calculated for a particular size $n$ of the truncation, yield weaker estimates, than the LTI SDP of the same size. Concretely, we show that the finite size criteria bounds lie in the feasible set of the LTI SDP. The above relation relies on finding a positive semidefinite generating term $q_{n}(\delta)$ for the truncated operators $Q_{n}(\delta)$ corresponding to each of the bounds. We show the inclusion for the martingale bound using the following positive local decomposition,
$\begin{split}q_{3l}(1-2\eta_{l}):=\eta_{l}(P_{\mathcal{I},\mathcal{I}+1}+P_{\mathcal{I}+1,\mathcal{I}+2})\\ +\{P_{\mathcal{I},\mathcal{I}+1},P_{\mathcal{I}+1,\mathcal{I}+2}\},\end{split}$ (35)
which assumes evaluating the LTI SDP for the coarse-grained Hamiltonian $K$ (as defined in Method 2 (section 3.1)). Similarly, the expression for the Knabe bound is,
@@
The detailed proofs of these claims as well as of the attainment of the Gosset–Mozgunov bound are given in Appendix B.
An important implication of eq. 34 is that the LTI-SDP does not necessarily lower bound the gap of the open boundary system, i.e. $\epsilon_{n}$, in the thermodynamic limit. This is because the formulation of $\delta_{\mathrm{LTI}}(n)$ only ensures that it lower bounds the gap of a periodic ring of size $m>2n$. This becomes evident by taking the $n\to\infty$ limit on both sides of eq. 34,
@@
4 Numerical benchmarking
In this section we present the results of the numerical application of the method presented in Section 2 to 1-dimensional spin-chain models. We compared our method with existing finite-size criteria, namely, the Knabe bound (eq. 23) and the bound due to Gosset and Mozgunov (eq. 24).
Because of the inclusions shown in Sec. 3, we know that for the same $n$ our method always gives an equal or tighter lower bound on the gap than either of those existing bounds. However, when comparing the methods we should keep in mind that the computational tasks involved are different. While the finite-size methods require the computation of the gap of the $n$-body OBC Hamiltonian—a task with a memory scaling of $\mathcal{O}(d^{n})$—our method optimizes over a positive semidefinite matrix of size $d^{n}\times d^{n}$. The results in this section show that our method outperforms the existing ones even when compared on equal footing (in terms of size of the largest variable in the problem): For the models on which we benchmarked, we observed that $\delta_{\mathrm{LTI}}(n)\geq\delta_{\mathrm{Knabe}}(2n)$, and similarly for $\delta_{\mathrm{GM}}(2n)$.
@@ 4 Numerical benchmarking
Before presenting the results we address a technical but important issue: When performing the optimization on a computer the SDP solver will produce a solution which is feasible up to some finite precision. If we wish to regard such a solution as providing a mathematical proof of the existence of a gap, we have to make sure that we have a way to obtain a strictly feasible solution from the one produced by the solver. In our case we obtain a numerical solution for the variable $q_{n}$ which should serve as our certificate of a gap (via equation eq. 7). In practice we obtain $q_{n}\succeq-\epsilon$ where $\epsilon>0$ is the solver precision. For the LTI SDP one cannot simply add $\epsilon\mathbb{I}$ to $q_{n}$ as this would shift the global energy by an extensive amount and a more refined approach has to be taken. In Appendix D, we explain how we modified the LTI SDP to obtain certified solutions. When testing the method on gapless frustration free models and ones with very small gaps, we observed that the correction detailed in Appendix D is necessary. Without it the solver may produce wrong results (i.e. detect a finite gap when the model is gapless or detect a gap larger than the true one).
4.1 The AKLT model
@@
The AKLT model was the first example of a provably gapped isotropic spin chain with integral spin [2]. When the model was initially introduced it was proven to be gapped but an explicit lower bound was not provided [3]. In the same year, Knabe introuced his method and provided a lower bound of $0.24806$ for the gap of the AKLT model. Garcia-Saez et al. [22], using a method based on the rotation symmetry of the Hamiltonian, estimated the AKLT gap to be $0.350$, which is compatible with both exact diagonalization results for finite periodic chains $(0.35012$ for $n=16$ [22]), as well as DMRG calculations for finite open chains ($0.35037$ for $n=100$  [^2] Mingru Yang, private communication). Our method proves a lower bound of $\Delta_{\text{AKLT}}\geq 0.34976$ which was obtained by solving the LTI SDP eq. 11 for $n=6$.
@@
Figure 2: The gap of the $S=1$ AKLT chain. Lower bounds on the energy gap above the ground state of the infinite AKLT chain obtained with different methods are plotted as a function of the size $n$ of the finite system: The Knabe bound eq. 23, the Gosset–Mozgunov bound eq. 24, and our LTI SDP method eq. 11. In addition, gaps of finite systems with periodic boundary conditions (PBC) obtained with exact diagonalization (ED) are plotted for rings of sizes $n=8,12$, and $16$ by dotted, dashed, and solid horizontal black lines. Note that those lines lie essentially on top of each other in the plot (the value for $n=12$ and $n=16$ differs in the 6th digit). The LTI SDP significantly outperforms the other two methods and gives a lower bound of $0.34976$ for the AKLT gap.
Figure 2 shows the Knabe (eq. 23) and Gosset–Mozgunov (eq. 24) lower bounds, and the lower bounds computed with our method (the solution of eq. 11), all for different values of $n$. In addition the finite system gaps obtained using exact diagonalization for periodic boundary conditions (PBC) are plotted for $n=8,12$ and $16$ as horizontal lines. Note that the exact diagonalization results for the PBC chain are neither lower nor upper bounds on the gap in the thermodynamic limit. They are plotted for reference because the Knabe and Gosset–Mozgunov bounds, as well as the LTI SDP all provide lower bounds to the finite periodic systems’ gaps as well.
We observe that while all three methods (ours, Knabe, and Gosset–Mozgunov) detect the gap starting from system size $n=4$, the LTI SDP performs significantly better than any of the other methods. This is true even if we compare $\delta_{\mathrm{LTI}}(n)$ with $\delta_{\mathrm{Knabe}}(2n)$ and $\delta_{\mathrm{GM}}(2n)$. Recall that the LTI SDP eq. 11 provides a lower bound on the gap of any periodic system of size $\geq 2n$. Our results for the AKLT model show that the LTI SDP gets very close to this optimal value: for $n=6$ we have $\Delta_{\text{PBC}}(12)-\delta_{\text{LTI}}(6)=0.00036$.
4.2 Deformed clock models
@@
Due to the fact that the deformation preserves the product-state structure of the ground states of eq. 38, the martingale method can be used to prove that the model is gapped in the whole parameter range [44]: the overlaps between the the ground-state projectors are easy to compute analytically and $\eta$ in eq. 31 is thus shown to decay with the size of the support of the state ($l$ in eq. 31) (see also [53] for a specific treatment of this family of models). This strategy for proving the existence of a gap, however, has the disadvantage that it does not provide us with a concrete lower bound. The exact lower bound proven in this way has a prefactor: the gap of the blocked interaction term $\tilde{h}$ which acts on $2l\gg 1$ spins.
@@
We computed lower bounds on the gap for values of $(r,s)$ in the region $(0,1]\times(0,1]$ using the Knabe method (eq. 23), Gosset–Mozgunov (eq. 24) and our method (eq. 11). Figure 3 shows the regions in which each of the methods detected a gap. In the figure, the green-shaded region with dashed boundary indicates where Knabe’s criterion could detect a gap for system sizes $n=3,\ldots,10$, i.e., where $\max_{n\leq 10}\delta_{\mathrm{Knabe}}(n)>0$; the blue-shaded region with dash-dotted boundary marks the same for the Gosset–Mozgunov bound. The results obtained with our LTI method (eq. 11) are plotted in Figure 3 as individual data points. For each point on a grid in the $(r,s)$ plane we plotted a marker indicating the lowest $n$ for which a positive gap was detected by the LTI SDP (triangle for $n=3$, square for 4, and pentagram for 5).
Figure 3: Gaps of the $\mathbb{Z}_{3}$ family of deformed Potts clock models. For a family of spin chain models parameterized by two parameters $r,s$ (eq. 39) lower bounds on the gap were computed at different values of the parameters using different methods: the Knabe and Gosset–Mozgunov finite size criteria (eq. 23 and eq. 24 respectively), and our LTI SDP method eq. 11. The shaded regions indicate the parameter values for which the finite-size methods detected a gap using exact diagonalization results for systems up to size $n=10$ (green region with dashed border for Knabe and blue region with dash-dotted border for Gosset-Mozgunov). The different markers indicate parameter values for which the LTI SDP detected a gap, with different markers corresponding to different values of $n$ (see legend). For each coordinate $(r,s)$, we only plot the marker corresponding to the lowest value $n$ for which the LTI SDP detected a gap. Darker marker colors indicate larger gaps (see color bar). Each marker corresponds to the larger of the two gaps resulting from using either the original Hamiltonian or the modified Hamiltonian—in which the interaction term is a projection—as the input for the LTI SDP.
@@
4.3 1D Glauber model
@@
[…]
where $H(\sigma)$ represents a class of classical kinetic Ising models. For this class of Hamiltonians, the state above is found to be representable as a matrix product state, thus allowing for the construction of a parent Hamiltonian, i.e. a frustration-free Hamiltonian which has the state as its ground state [16]. Here we consider the family of parent Hamiltonians known as the 1D Glauber model [8], which has the local term
@@
i.e., independent of $\gamma$. Note that for calculating finite size criteria bounds, the local gap threshold $\epsilon_{n}$ (Sec. 3) corresponds to the gap of the projector Hamiltonian, which is independent of $\delta$.
Finally, to compute the finite size criteria bounds for a 3-local Hamiltonian, we have to adapt the calculation of $\delta_{\mathrm{Knabe}}$ and $\delta_{\mathrm{GM}}$ using a two-step coarse-graining procedure. We first block two $\mathrm{spin}-\frac{1}{2}$ particles of the original chain into one $\mathrm{spin}-\frac{3}{2}$ $(d=4)$ particle,
$\mathcal{I}=\{2i-1,2i\},$ (47)
@@
$\tilde{h}_{\mathcal{I},\mathcal{I}+1}$ $=h_{2i-1,2i,2i+1}+h_{2i,2i+1,2i+2},$ (48)
@@
Figure 4: Gaps of the 1D Glauber Hamiltonian family. For a family of spin chain models parameterized by two parameters $\gamma,\delta$ (eq. 44) lower bounds on the gap were computed at different values of the parameters using different methods: the Knabe and Gosset–Mozgunov finite size criteria (eq. 23 and eq. 24 respectively) computed for the coarse-grained Hamiltonian $\tilde{h}$ (eq. 48), and our LTI SDP method eq. 11 computed for the original Hamiltonian. The green shaded region (dashed border) marks the region in which the Knabe method detected a gap ($\gamma\lesssim 0.92$), the blue shaded region (dash-dotted border) marks the region where the Gosset–Mozgunov bound did ($\gamma\lesssim 0.96$). The gaps detected by our LTI method are plotted as individual data points with marker colors corresponding to the size of the gap (see color bar).
@@ 5 Conclusion and Outlook
An issue we did not resolve in this paper is the completeness of our hierarchy. Therefore, it remains open whether the hierarchy always detects a gap at some finite level when a model is indeed gapped, and whether the certified gap always converges to the true gap as the level in the hierarchy is increased. In this context, it is important to note the undecidability of the spectral gap problem for frustration-free Hamiltonians on 2D lattices [18, 19]. Coupling a complete hierarchy for lower bounds with a hierarchy of upper bounds obtained using variational methods would constitute a decision procedure for the spectral gap, violating the established undecidability results. We should therefore not expect to be able to prove completeness of our hierarchy in two or more spatial dimensions without further assumptions. Since the known 1D undecidability constructions [11] are not frustration-free, the question of completeness of the hierarchy for one-dimensional frustration-free systems remains open.
@@
Finally, from its formulation, we believe that our method can allow the certification of gaps for systems in higher spatial dimensions. The generalization of our method to such systems is straightforward. The challenge in applying our method in this setting lies in the ability to solve large scale SDPs. For a matrix variable of a given size, the complexity solving an SDP exceeds that of full exact diagonalization, which is required in order to certify the existence of a gap using finite size criteria. Yet, from the results we presented for one-dimensional systems it seems reasonable to expect that the system size for which one would need to solve the problem could be significantly smaller than the system sizes involved in proving gaps using finite size criteria in 2D (e.g. [39, 50, 38]).
We thank Milán Rozmán for guiding us through the martingale literature. We thank Mingru Yang for enlightening discussions as well as for providing us with the DMRG result for the AKLT model. We thank Ignacio Cirac for stimulating discussions at the beginning of the project. We thank Jin-Fu Chen for useful discussions about the Glauber model. We thank Vedran Dunjko for useful discussions about undecidability. We thank Anurag Anshu for bringing the coarse-graining method for higher-locality Hamiltonians to our attention. I.K. and N.S. acknowledge support from the European Union’s Horizon 2020 research and innovation programme through Grant No. 863476 (ERC-CoG SEQUAM), and the FWF Quantum Austria Funding Initiative project “Entanglement Order Parameters” (Grant DOI 10.55776/P36305), funded through the European Union – NextGenerationEU. N.S. also acknowledges support by the the Austrian Science Fund FWF through Grant DOIs 10.55776/COE1 and 10.55776/F71. The computational results presented have been achieved in part using the Austrian Scientific Computing (ASC) infrastructure. K.S.R. and J.T. acknowledge the support received from the European Union’s Horizon Europe research and innovation programme through the ERC StG FINE-TEA-SQUAD (Grant No. 101040729). P.E. and J.T. acknowledge the support received by the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme. P.E. acknowledges the support received through the NWO-Quantum Technology programme (Grant No. NGF.1623.23.006). P.E. further acknowledges funding by the Carl-Zeiss-Stiftung (CZS Center QPhoton). This publication is part of the ‘Quantum Inspire - the Dutch Quantum Computer in the Cloud’ project (with project number [NWA.1292.19.194]) of the NWA research program ‘Research on Routes by Consortia (ORC)’, which is funded by the Netherlands Organization for Scientific Research (NWO). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the funding institutions. Neither of the funding institutions can be held responsible for them. Parts of this work were performed by using the compute resources from the Academic Leiden Interdisciplinary Cluster Environment (ALICE) provided by Leiden University. F.B. acknowledges financial support from the Alexander von Humboldt Foundation and support from the ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union – NextGenerationEU.